Two airplanes leave the Atalnta airport at th esame time, traveling in opposite directions. One plane travels 30 miles per hour faster than the other. After 3 hours, the planes are 3150 miles apart. what is the rate of each plane?
Answer by checkley71(8403)

3150=3*X+3(X+30)
3150=3X+3X+90
6X=3150-90
6X=3060
X=3060/6
X=510 MPGH FOR THE SLOWER PLANE &
510+30=540 MPH FOR THE FASTER PLANE PROOF
3150=3+*510+3*540
3150=1530+1620
3150=3150